Are your finances, buying habits, medical records, and phone calls really private? A real concern for many adults is that computers and the Internet are reducing privacy. A survey conducted by Peter D. Hart Research Associates for the Shell Poll was reported in USA Today. According to the survey, 47% of adults are concerned that employers are monitoring phone calls. Use the binomial distribution formula to calculate the probability of the following.

Out of five adults, exactly three are concerned that employers are monitoring phone calls. (Round your answer to three decimal places.)
Answer #1

Here, n = 5, p = 0.47, (1 - p) = 0.53 and x = 3
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(X = 3)
P(X = 3) = 5C3 * 0.47^3 * 0.53^2
P(X = 3) = 0.292

ans: 0.292